OpenStudy (anonymous):

what is the product of t sqrt (8-3 sqrt5)

OpenStudy (anonymous):

\[\sqrt{t}(8-3\sqrt{5})\]

OpenStudy (gorv):

is it 7

OpenStudy (anonymous):

no a t

OpenStudy (anonymous):

@gorv

OpenStudy (gorv):

calculate sqrt 5 first

OpenStudy (anonymous):

\[8\sqrt{t}-3\sqrt{5t}\]

OpenStudy (anonymous):

that what i got for an answer

OpenStudy (gorv):

yeah but u can calculate sqrt 5

OpenStudy (gorv):

sqrt5=2.23 3*2.23sqrt(t)=4.47sqrt(t)

OpenStudy (gorv):

8sqrt(t)-4.47sqrt(t)=3.53sqrt(t)

OpenStudy (gorv):

@allie